Allocation: Fixed and Variable Costs, Budgeted Fixed and Variable Costs

Biotechtron, Inc., has two research laboratories in the Southwest, one in Yuma, Arizona, and the other in Bernalillo, New Mexico. The owner of Biotechtron centralized the legal services function in the Yuma office and had both laboratories send any legal questions or issues to the Yuma office. The legal services support center has budgeted fixed costs of $151,000 per year and a budgeted variable rate of $60 per hour of professional time. The normal usage of the legal services center is 2,400 hours per year for the Yuma office and 1,600 hours per year for the Bernalillo office. This corresponds to the expected usage for the coming year.

Required:

1. Determine the amount of legal services support center costs that should be assigned to each office. In your computations, carry ratio values out to two decimal places. Round your final answers to the nearest dollar if rounding is required.

Yuma $
Bernalillo $

2. Since the offices produce services, not tangible products, what purpose is served by allocating the budgeted costs?

3. Now, assume that during the year, the legal services center incurred actual fixed costs of $154,100 and actual variable costs of $245,500. It delivered 4,070 hours of professional time—2,830 hours to Yuma and 1,240 hours to Bernalillo.

Determine the amount of the legal services center's costs that should be allocated to each office.

Yuma $
Bernalillo $

4. Did the costs allocated differ from the costs incurred by the legal services center?

If so, why?
Answer #1
1 Yuma Bernalillo
Ratio for fixed costs* 0.6 0.4
Fixed costs 90600 60400
Variable costs** 144000 96000
234600 156400
*Yuma = 2,400/4,000 = 0.60; Bernalillo = 1,600/4,000 = 0.40
**Yuma = $60 × 2,400; Bernalillo = $60 × 1,600
2 Costing out services serves the same purposes as costing out tangible products (e.g., pricing, profitability analysis, and performance evaluation).Once the costs are allocated to each revenue-producing center, then the costs must be assigned to individual services through the use of an overhead rate or rates
3 If the purpose is to cost out individual services, then the allocation is identical to that given in Requirement 1.If the purpose is for performance evaluation, then variable costs equal the predetermined rate multiplied by the actual usage. The fixed costs are allocated the same way as before
Yuma Bernalillo
Variable costs** 169800 74400
Fixed costs 90600 60400
260400 134800
**Yuma = $60 × 2,830; Bernalillo = $60 × 1,240
4 The allocated costs of $395,200 were $4,400 lower than the actual costs of$399,600, because the producing departments are charged an allocation based on budgeted costs rather than actual costs. Budgeted costs are allocated so that the efficiencies or inefficiencies of the service center are not assigned to the user departments.